OpenStudy (anonymous):

How is this simplified, really confused. (x-1)^2-(x-1) let f(x)=x^2-x and g(x)=x-1. Find (f o g)(x) and (g o f)(x) so I got, (f o g)(x)=f(g(x))=f(x-1)=(x-1)^2-(x-1), the answer says its, x^2-2x+1-x+1 so (f o g)(x)=x^2-3x+2

OpenStudy (anonymous):

I cam up with this; x^2-x+1-2x

OpenStudy (trantom):

you want to use the distributive property first (x-1)^2-(x-1) x^2-1-(x-1) For this step the 1 that isn't in parenthesis is actually 1^2 but 1^2 is 1 so I left it at that Then you get rid of parenthesis You will notice I make the (x-1) into x+1. This is because if you subtract a difference you end up adding the second term when you get rid of parenthesis x^2-1-x+1 Combine like terms Since -1+1 is 0, you can get rid of constants altogether. x^2-x The final answer is x^2-x although I should probably say "the simplified version of the expression (x-1)^2-(x-1) is x^2-x" because that is a lot more clear xD

OpenStudy (trantom):

So yea the answer is x^2-x

OpenStudy (anonymous):

I have the answer in the book, it says x^2-2x+1-x+1=x^2-3x+2

OpenStudy (trantom):

really?

OpenStudy (trantom):

by the way please don't pay attention to mail from me malibug

OpenStudy (trantom):

why is there an equal sign in the answer? the question doesn't have one...

OpenStudy (anonymous):

Ok I will put the whole question up because I think this confusing. I only the one part up because I was only stuck on that part.

OpenStudy (trantom):

oh what! yea always put up the whole thing

OpenStudy (anonymous):

Sorry about that. I am confused how they got the simplified part?

myininaya (myininaya):

\[(x-1)^2=(x-1)(x-1) \text{ \not } x^2-1\]

myininaya (myininaya):

\[(x-1)^2=(x-1)(x-1)=x(x-1)-1(x-1)=x^2-x-x+1=x^2-2x+1\] So you started out with: \[(x-1)^2-(x-1)=(x^2-2x+1)-x+1\] You can combine like terms from here

OpenStudy (anonymous):

@malibugranprix2000, do you still need help with this?

OpenStudy (anonymous):

@GABI.J I actually figured it out. Thanks.

OpenStudy (anonymous):

Anytime, as long as you understand....

OpenStudy (anonymous):

Thanks @myininaya really helped.